# RIDE: Difficulty Evolving Perturbation with Item Response Theory for Mathematical Reasoning

### v1

@@ 4.2 Reinforcement Learning for Difficulty Evolution
We use Qwen3-8B (qwen3technicalreport) as the base model and train a math question rewriting model via Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). The SFT stage primarily serves as a cold start, which equips the base model with an initial ability to rewrite original math questions $q$ into more difficult ones $q^{\prime}$ and standardizes the output format. We curate $2,000$ high-quality rewritten questions $\mathcal{D}$ by distilling and filtering GPT-5-mini  [^2] https://openai.com/ outputs, and used this corpus to perform full-parameter fine-tuning of Qwen3-8B:
@@
$\theta$ denotes to the original parameter of the base model and $\theta_{SFT}$ denotes to the new parameter after full-parameter fine-tuning.
In the RL stage, we use the Group Sequence Policy Optimization (GSPO) algorithm (zheng2025groupsequencepolicyoptimization) for optimization. Starting from SFT model with parameter $\theta_{\text{SFT}}$, for each $q$ we sample a group $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}\sim\mathcal{P}_{\theta}(\cdot|q)$ with group size $K$, compute rewards $R(q,q_{k}^{\prime})$ and group-normalized advantages, and update model parameter $\theta$ with GSPO. Our rewards consist of:
@@
  • Keyword and Length Reward. To enforce a parsable format and penalize overly verbose rewrites, we add auxiliary keyword and length rewards $R_{\text{key}}$ with small coefficients to avoid overshadowing difficulty and correctness rewards.
Given reward weights $\alpha$, $\beta$, and $\gamma$, the final reward $R$ is defined as their weighted sum. For each $q$ with samples $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}$, let $R_{k}\equiv R(q,q_{k}^{\prime})$ and compute advantage $A_{k}$:
$R=\alpha R_{\text{diff}}+\beta R_{\text{cor}}+\gamma R_{\text{key}},\qquad A_{k}=\frac{R_{k}-\tfrac{1}{K}\sum_{k=1}^{K}R_{k}}{\mathrm{std}(R_{1:K})},$
@@
${\rho}_{k}(\theta)=\left(\frac{\mathcal{P}_{\theta}(q_{k}^{\prime}\mid q)}{\mathcal{P}_{\theta_{\text{old}}}(q_{k}^{\prime}\mid q)}\right)^{\!1/|q_{k}^{\prime}|}=\exp\!\left(\frac{1}{|q_{k}^{\prime}|}\sum_{t=1}^{|q_{k}^{\prime}|}\log\frac{\pi_{\theta}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}{\pi_{\theta_{\text{old}}}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}\right)$
@@
$\mathcal{L}_{\mathrm{GSPO}}(\theta)=\mathbb{E}_{q\in\mathcal{D}}\Bigg[\frac{1}{K}\sum_{k=1}^{K}\min\!\big(\rho_{k}(\theta)\,A_{k},\ \mathrm{clip}(\rho_{k}(\theta),\,1-\varepsilon,\,1+\varepsilon)\,A_{k}\big)\Bigg]$
@@
Once the reward converges, we obtain our final difficulty-aware math rewriting model RIDE-8B.
Considering that existing mathematical datasets may not provide complete solutions, we only take the original question $q_{\text{original}}$ and its answer $a_{\text{original}}$ from benchmarks or training data as input. The corresponding prompt is shown in Appendix B.3. Instead of prescribing specific rules, the prompt provides high-level guidance for increasing difficulty, allowing the model to freely rewrite questions, thereby creating unpredictable paths of difficulty progression. Subsequently, we take the prompt context as input into RIDE-8B, which infers the rewritten question $q_{\text{rewrite}}$, solution $s_{\text{rewrite}}$, and answer $a_{\text{rewrite}}$:
@@ 5.1 Experiment Setup
We use our rewriting model RIDE-8B to perform multi-round difficulty evolution on the competition-level mathematical-reasoning benchmarks AMC-23 and AIME-24, which contain $40$ and $30$ questions, respectively. Using GPT-5, we filter out questions that are unsolvable or have incorrect answers to construct the final perturbed benchmarks RIDE-AMC and RIDE-AIME, which are further manually corrected to ensure answer accuracy, with annotators holding at least a bachelor’s degree. We evaluate the pass@1 performance of cutting-edge open-source and proprietary models in the zero-shot setting. In addition, we extract data from the DeepMath dataset for rewriting, and after filtering with DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport), we obtain 1010K high-quality mathematical reasoning samples, referred to as RIDE-DeepMath, which serve as training data to improve the models’ ability for mathematical reasoning.
5.2 Main Results
We evaluate $23$ LLMs comprising $8$ proprietary models (including GPT-5 and Grok-4) and $15$ open-source models (including DeepSeek-V3.1 and Kimi-K2-Instruct (kimiteam2025kimik2openagentic)), spanning parameter counts from 0.60.6B to 11TB and covering multiple model series. We report pass@1 performance on the original AMC-23 and AIME-24 benchmarks, and then apply the same prompts to our perturbed benchmarks, RIDE-AMC and RIDE-AIME, again recording pass@1 performance. To quantify robustness, we compute the performance drop rate (PDR) (li2024gsmpluscomprehensivebenchmarkevaluating):
where $x$ and $y$ denote the question and the corresponding answer, respectively, $\mathcal{D}$ is the original benchmark (AMC-23 or AIME-24), $\mathcal{D}_{a}$ is the corresponding perturbed benchmark (RIDE-AMC or RIDE-AIME), $\mathrm{LM}(\cdot)$ is the model’s prediction, and $\mathbb{I}[\cdot]$ is the indicator function. Lower PDR indicates greater robustness. Results are presented in Table 5.2.
@@ 5.2 Main Results
5.3 Case Study
Figure 3 illustrates the rewritten question. The original question is a classic coupon collector question with replacement. The rewritten question (1) clarifies the “pick 10 marbles at a time with replacement” procedure without changing the core idea; (2) raises the target success probability from 9595% to 9999%, increasing the required trials; and (3) requires using the Poisson approximation and the union bound, increasing restrictions and shifting the solution from exact combinatorial counting to multi-step reasoning. From this example, we can see that without explicitly stating a specific rule, our method rewrites the question by modifying the wording, values, and conditions. This raises the difficulty while maintaining the same core concept being tested.
Figure 3: Case study for our method. The original question is sampled from DeepMath dataset, and rewritten question is the corresponding rewritten version in RIDE-DeepMath.
5.4 Analysis
In this subsection, we will conduct more experiments to further explore and analyze the effectiveness and impact of our method.
Method Dataset Average Consistent
Rule-based AMC-23 40.83 28.33
RIDE AMC-23 58.75 48.33
Rule-based AIME-24 36.67 28.89
RIDE AIME-24 63.33 55.56
Table 2: Win rate comparison of the rule-based method and RIDE. Average win rate measures selections across both judgments, while consistent win rate measures questions won in both judgments.
Win Rate Comparison of RIDE and Rule-based Method. Following GSM-Plus (li2024gsmpluscomprehensivebenchmarkevaluating), we use Qwen3-8B to generate perturbed AMC-23 data with the same prompting strategy. To evaluate rewrite quality, we design prompts assessing completeness, clarity, conceptual consistency, structural perturbations, and avoidance of shortcuts (Appendix B.3). GPT-5 acts as an evaluator to compare rule-perturbed questions with RIDE-AMC and judge each pair twice with swapped order to mitigate position bias. We compute two win rate metrics: (1) Average win rate—the proportion of times a method is selected across both evaluations; (2) Consistent win rate—the proportion of questions where a method is selected in both evaluations. We run three rounds and report average results. As shown in Table 2, our method significantly outperforms the rule-based approach on both metrics, demonstrating that RIDE rewrites achieve higher quality.
Figure 4: Comparison of pass@n performance before and after benchmark perturbation.
@@ 5.4 Analysis
Data Augmentation For Training Data. Our method also serves as an effective data augmentation strategy for mathematical reasoning. Using RIDE-8B, we applied difficulty-evolution rewriting to the DeepMath dataset and filtered incorrect samples with DeepSeek-V3.1, yielding RIDE-DeepMath with over 10K examples. Without SFT, we trained Qwen3-0.6B and Gemma3-4B-It via reinforcement learning with GSPO, producing RIDE-Qwen-0.6B and RIDE-Gemma-4B. Results (Table 4) show that models trained with reinforcement learning on our data achieve improvements over their base models. In particular, RIDE-Qwen-0.6B and RIDE-Gemma3-4B have surpassed GPT-4o and Qwen2.5-Math-72B, and are close to Qwen3-4B in the non-thinking mode.
@@
Model Method Data Augmentation AUC-ROC
VAE Sampling
1-PL VI - - 89.81
1-PL VI - 91.01
1-PL VI - 91.06
[…]
1-PL MCMC - - 89.81
2-PL VI - - 85.18
@@
Table 4: Comparison of different IRT models and methods. VI denotes Variational Inference and $MC$ denotes Markov Chain Monte Carlo. We select the method combination with the highest AUC-ROC as IRT parameter estimation approach.
Effect of IRT models. We compare different IRT models and data augmentation methods for question difficulty estimation. The 2-PL model adds a discrimination parameter to the 1-PL, while the 3-PL further introduces a guessing parameter. We randomly mask 20% of the response matrix for augmentation and model fitting, and evaluated performance using AUC-ROC. As shown in Table 4, augmentation generally improves prediction accuracy, confirming its effectiveness. Given the high cost of MCMC, we ultimately adopt the 1-PL model with both augmentation methods and VI for parameter estimation, achieving a balance between accuracy and efficiency. Detailed results and the effect of the difficulty ranker are provided in Appendix C.
Statistics of Different Rewriting Models. We train RIDE-1.7B/4B based on Qwen3-1.7B/4B and keep SFT checkpoints (RIDE-1.7B-SFT, RIDE-4B-SFT, RIDE-8B-SFT) to disentangle SFT from RL effects. The questions rewritten by each model are evaluated in terms of generation time, token count, correctness, and semantic similarity. Results (Figure 5) show that compared with Qwen3-8B, both SFT and RL models produce longer outputs and run slower, indicating that rewriting expands question statements and solutions. Semantic similarity drops slightly, reflecting greater diversity while remaining close to the original questions. RL improves correctness over both the base and SFT-only models and RIDE-8B/4B achieve markedly higher correctness than RIDE-1.7B, suggesting that parameter scale benefits correctness, while the difference between 4B and 8B is small, implying limited scale effects in this range.
Figure 5: Statics of average token count, average inference time, correct rate and cosine similarity across different models.

### v2

@@ 4.2 Reinforcement Learning for Difficulty Evolution
We use Qwen3-8B (Team, 2025) as the base model and train a math question rewriting model via Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). The SFT stage primarily serves as a cold start, which equips the base model with an initial ability to rewrite original math questions $q$ into more difficult ones $q^{\prime}$ and standardizes the output format. We curate $2,000$ high-quality rewritten questions $\mathcal{D}$ by distilling and filtering GPT-5-mini  [^2] https://openai.com/ outputs, and used this corpus to perform full-parameter fine-tuning of Qwen3-8B:
@@
$\theta$ and $\theta_{SFT}$ denote to the original parameter of the base model and the new parameter after full-parameter fine-tuning, respectively.
In the RL stage, we use the Group Sequence Policy Optimization (GSPO) algorithm (Zheng et al., 2025) for optimization. Starting from SFT model with parameter $\theta_{\text{SFT}}$, for each $q$ we sample a group $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}\sim\mathcal{P}_{\theta}(\cdot|q)$ with group size $K$, compute rewards $R(q,q_{k}^{\prime})$ and group-normalized advantages, and update model parameter $\theta$ with GSPO. Our rewards consist of:
@@
Given reward weights $\alpha$ and $\beta$, the final reward $R$ is defined as their weighted sum. For each $q$ with samples $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}$, let $R_{k}\equiv R(q,q_{k}^{\prime})$ and compute advantage $A_{k}$:
$R=\alpha R_{\text{diff}}+\beta R_{\text{cor}},\quad A_{k}=\frac{R_{k}-\tfrac{1}{K}\sum_{k=1}^{K}R_{k}}{\mathrm{std}(R_{1:K})},$
@@
${\rho}_{k}(\theta)$ $=\left(\frac{\mathcal{P}_{\theta}(q_{k}^{\prime}\mid q)}{\mathcal{P}_{\theta_{\text{old}}}(q_{k}^{\prime}\mid q)}\right)^{\!1/|q_{k}^{\prime}|}$
$=\exp\!\left(\frac{1}{|q_{k}^{\prime}|}\sum_{t=1}^{|q_{k}^{\prime}|}\log\frac{\pi_{\theta}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}{\pi_{\theta_{\text{old}}}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}\right)$
@@
$\mathcal{L}_{\mathrm{GSPO}}(\theta)=$ $\mathbb{E}_{q\in\mathcal{D}}\Bigg[\frac{1}{K}\sum_{k=1}^{K}\min\Big(\rho_{k}(\theta)A_{k},$
$\text{clip}\big(\rho_{k}(\theta),1-\varepsilon,1+\varepsilon\big)A_{k}\Big)\Bigg]$
@@
Once the reward converges, we obtain difficulty-aware math rewriting model RIDE-8B. Considering that existing mathematical datasets may not provide complete solutions, we only take the original question $q_{\text{original}}$ and its answer $a_{\text{original}}$ from benchmarks or training data as input. The corresponding prompt is shown in Appendix A.3. Instead of prescribing specific rules, the prompt provides high-level guidance for increasing difficulty, allowing the model to freely rewrite questions, thereby creating unpredictable paths of difficulty progression. Subsequently, we take the prompt context as input into RIDE-8B, which infers the rewritten question $q_{\text{rewrite}}$, solution $s_{\text{rewrite}}$, and answer $a_{\text{rewrite}}$:
@@ 5.1 Experiment Setup
We use our rewriting model RIDE-8B to perform multi-round difficulty evolution on the competition-level mathematical-reasoning benchmarks AMC-23 and AIME-24, which contain $40$ and $30$ questions, respectively. Using GPT-5, we filter out questions that are unsolvable or have incorrect answers to construct the final perturbed benchmarks RIDE-AMC and RIDE-AIME, which are further manually checked to ensure answer accuracy, with annotators holding at least a bachelor’s degree. We evaluate the pass@1 performance of cutting-edge open-source and proprietary models in the zero-shot setting. In addition, we extract data from the DeepMath dataset for rewriting, and after filtering with DeepSeek-V3.1 (DeepSeek-AI, 2024), we obtain 1010K high-quality mathematical reasoning samples, referred to as RIDE-DeepMath, which serve as training data to improve the models’ ability for mathematical reasoning.
@@
5.2 Main Results
We evaluate $23$ LLMs comprising $8$ proprietary models (including GPT-5 and Grok-4) and $15$ open-source models (including DeepSeek-V3.1 and Kimi-K2-Instruct (Team et al., 2025)), spanning parameter counts from 0.60.6B to 11TB and covering multiple model series. We report pass@1 performance on the original AMC-23 and AIME-24 benchmarks, and then apply the same prompts to our perturbed benchmarks, RIDE-AMC and RIDE-AIME, again recording pass@1 performance. To quantify robustness, we compute the performance drop rate (PDR) (Li et al., 2024):
where $x$ and $y$ denote the question and the corresponding answer, respectively, $\mathcal{D}$ is the original benchmark (AMC-23 or AIME-24), $\mathcal{D}_{a}$ is the corresponding perturbed benchmark (RIDE-AMC or RIDE-AIME), $\mathrm{LM}(\cdot)$ is the model’s prediction, and $\mathbb{I}[\cdot]$ is the indicator function. Lower PDR indicates greater robustness. Results are presented in Table 1.
@@ 5.2 Main Results
6 Analysis
In this subsection, we conduct more experiments to explore and analyze the effectiveness and impact of our method.
@@ 6 Analysis
Figure 3: Comparison of pass@n performance before and after benchmark perturbation. Despite scaling improvements with $n$, the persistent performance gap confirms RIDE’s stability.
Win Rate Comparison of RIDE and Other methods. To evaluate rewrite quality, we design prompts assessing completeness, clarity, conceptual consistency, structural perturbations, and avoidance of shortcuts (Appendix A.3). GPT-5 acts as an evaluator to compare the outputs from different methods. Each pair is judged twice with swapped order to mitigate position bias. We compute two win rate metrics: (1) Average win rate—the proportion of times a method is selected across both evaluations; (2) Consistent win rate—the proportion of questions where a method is selected in both evaluations. Ties are allowed in the evaluator’s output. We run three rounds and report average results. We compare our method with the following method: (1) Rule-based method. Following GSM-Plus (Li et al., 2024), we use Qwen3-8B to generate perturbed AMC-23 and AIME-24 data with the same prompting strategy as rule-based baseline. (2) Contrastive prompting. Following the contrastive prompting method (Yao, 2025), we instruct Qwen3-8B to generate an easier and a more difficult rewrite at the same time and choose the difficult one as the final rewrite question. (3) SFT-only model and RL model without difficulty reward. We generate rewritten questions using the SFT-only model (without RL) and an RL model trained without the difficulty reward $R_{\text{diff}}$ as part of our ablation study. As shown in Table 2, our method significantly outperforms the rule-based and contrastive prompting on both metrics, demonstrating that RIDE rewrites achieve higher quality. Ablation study also shows the effect of our RL stage and difficulty reward.
Compared Method Dataset Average Consistent
Baseline RIDE Baseline RIDE
Rule-based AMC-23 40.83 58.75 28.33 48.33
AIME-24 36.67 63.33 28.89 55.56
Contrastive Prompting AMC-23 49.17 50.83 40.00 40.83
AIME-24 32.78 67.22 26.67 61.11
SFT-only AMC-23 38.75 60.00 33.33 53.33
AIME-24 36.67 62.22 26.67 52.22
w/o $R_{\text{diff}}$ AMC-23 43.75 54.17 35.83 47.50
AIME-24 27.78 72.22 21.11 65.56
Table 2: Win rate comparison of RIDE and other methods.
Case Study and Rewrite Strategy Analysis. Figure 4(a) illustrates the rewritten question. The original question is a classic coupon collector question with replacement. The rewritten question (1) clarifies the “pick 10 marbles at a time with replacement” procedure without changing the core idea; (2) raises the target success probability, increasing the required trials; and (3) requires using the Poisson approximation and the union bound, increasing restrictions. To systematically analyze these rewrites, we distill six core rewrite strategies from the questions in the RIDE-AIME and RIDE-AMC datasets: (1) Wording and Paraphrasing Modification, (2) Distracting Info, (3) Numerical Substitution, (4) Add Extra Steps, (5) Change Constraints, and (6) Change Target. (Detailed descriptions are provided in the Appendix A.4). Our analysis (Figure 4(b)) reveals that these strategies are frequently combined. Over 60% of the rewritten questions are composed of two or more strategies. To further investigate the impact on reasoning depth, we categorize strategies 11-33 as Surface Strategies, which have a minor impact on reasoning depth, and strategies 44-66 as Deep Strategies, which significantly affect the required reasoning. Statistics show that rewrites using a mix of both Surface and Deep strategies are the most common (44.3%). The accuracy of Qwen2.5-Math-72B and Llama4-Scout on questions corresponding to each strategy shows that performance varies across different strategies. In summary, our method can generate combinations of one or more rewrite strategies without explicitly defining specific rules. These combinations effectively provide a difficulty perturbation for the original question.
(a) Case study for our method. The original question is sampled from DeepMath dataset, and rewritten question is the corresponding rewritten version in RIDE-DeepMath.
(b) Rewrite strategy analysis.
Figure 4: Case study and analysis. (a) The original and rewritten questions from DeepMath and RIDE-DeepMath. (b) The analysis of rewrite strategies.
The Relation Between Difficulty and Robustness. To validate whether the performance drop stems from an increase in human-perceived difficulty or from the model’s inherent vulnerability, we conduct an analysis using the surface-level rewriting dataset mentioned above, which is assumed to keep human cognitive difficulty invariant. We record the transitions between “Correct on Original, Incorrect on Rewritten” and the inverse for each paired sample and apply McNemar’s test. The resulting p-value of $2.3847\times 10^{-11}$, being lower than $0.05$, reveals a significant asymmetry. This demonstrates that the performance drop is driven by the model’s non-robustness rather than shifts in difficulty. Additionally, as detailed in the Appendix B.1, we categorize questions into three difficulty levels and the consistent degradation across all difficulty levels (e.g., $11.43\%$ on ”Easy” questions and $12.86\%$ on ”Medium” questions) suggests this drop is independent of reasoning depth. From a theoretical perspective, the IRT-based difficulty is statistically estimated from the model’s response patterns, which is highly correlated with empirical accuracy. This difficulty focuses on the probability of a correct response rather than directly reflecting human cognitive complexity. Consequently, our “difficulty evolution” essentially transforms questions into formats that are challenging for the model to solve, thereby serving as a robustness test.
@@
Data Augmentation For Training Data. Our method also serves as an effective data augmentation strategy for mathematical reasoning. Using RIDE-8B, we applied difficulty-evolution rewriting to the DeepMath dataset and filtered incorrect samples with DeepSeek-V3.1, yielding RIDE-DeepMath with over 10K examples. Without SFT, we trained Qwen3-0.6B, Qwen3-1.7B and Gemma3-4B-It via reinforcement learning with GSPO, producing RIDE-Qwen-0.6B-Thinking, RIDE-Qwen-1.7B and RIDE-Gemma-4B. Results (Table 3) show that models trained with reinforcement learning on our data achieve improvements over their base models. In particular, RIDE-Qwen-0.6B-Thinking and RIDE-Gemma3-4B have surpassed GPT-4o and Qwen2.5-Math-72B, and are close to Qwen3-4B in the non-thinking mode.
@@
Effect Analysis. We additionally train RIDE-4B based on Qwen3-4B and keep SFT checkpoints (RIDE-4B-SFT, RIDE-8B-SFT) to disentangle SFT from RL effects. As shown in Table 4, Difficulty score is defined as the predicted probability from our difficulty ranker that the rewritten question is more difficult than the original. The experiments demonstrate that the RL-tuned RIDE-8B achieves the highest average Difficulty score, thereby validating the effect of our RL strategy. We also conduct ablation studies on the RIDE-4B model. We introduce two variations: (1) replacing the verifier for the correctness reward $R_{\text{cor}}$ with DeepSeek-V3.2 (RIDE-4B-DeepSeek), and (2) removing $R_{\text{cor}}$. We employ GPT-5 to evaluate the correctness of the rewritten questions generated by these models. The results indicate that RIDE-4B (GPT-5-mini verifier) achieves the highest correctness. RIDE-4B-DeepSeek model follows, as DeepSeek-V3.2 tends to output more false negatives (see Appendix B.4), whereas removing the correctness reward leads to decline in correctness. The DeepSeek verifier also improves correctness shows that our method is not heavily dependent on closed-source models, as open-source models are also effective. This confirms the validity and necessity of our correctness reward.
Besides, We compare different IRT models for question difficulty estimation. The 2-PL model adds a discrimination parameter to the 1-PL, while the 3-PL further introduces a guessing parameter. We randomly mask 20% of the response matrix for model fitting, and evaluate performance using AUC-ROC. We ultimately adapt the 1-PL model with VI for parameter estimation.
Part A: Effect analysis of RL and correctness reward
Model Difficulty Correctness
Qwen3-8B 74.61 40.00
RIDE-8B-SFT 78.78 40.00
RIDE-8B 80.18 42.50
RIDE-4B - 45.00
RIDE-4B-DeepSeek - 37.50
RIDE-4B w/o $R_{\text{cor}}$ - 32.50
Part B: Comparison of different IRT models
Model Method AUC-ROC
1-PL VI 91.29
1-PL MCMC 89.81
@@
Table 4: Effect analysis. Part A shows the ablation study on RL and correctness reward $R_{\text{cor}}$, while Part B compares different IRT models and parameter estimation methods.